# Kinematic cosmic dipole from a large sample of strong lenses

### 11 Mar 2026 v1

@@ Abstract
Measurements of the kinematic cosmic dipole continue to show an intriguing tension between the value inferred from the CMB and that obtained from high-redshift source number counts. While the measured dipole direction appears consistent, the amplitude, set by the observer’s peculiar velocity $v_{o}$, remains in significant disagreement. In this paper, we propose using strong gravitational lenses with well-measured Einstein radii to estimate the kinematic cosmic dipole, through the relativistic distortion of the Einstein angle induced by the observer’s motion. We show that this effect could be detected solely from measurements of the Einstein radius in wide, high-resolution imaging surveys such as Euclid. However, the precision achievable using Einstein-radius measurements alone, without redshift or lens-galaxy mass information, appears insufficient to discriminate between the CMB value of $v_{o}$ and that derived from source number counts at high statistical significance. Nevertheless, we demonstrate that including a large sample of lenses with available kinematic information, either via the Fundamental Plane relation or, ideally, through spectroscopic velocity-dispersion measurements, drastically reduces the noise and substantially improves the constraining power of this method. We show that, for a realistic sample of strong lenses detected by Euclid and complemented with spectroscopic velocity dispersion measurements from 4MOST or DESI, it is possible to discriminate between the CMB- and source-number-counts–inferred values at the $\sim 4\sigma$ level using a new, fully independent method. We further demonstrate that this technique is only weakly sensitive to strong-lensing selection effects, with selection biases and threshold effects estimated to be well below the 1% level.
@@ 1 Introduction
In this work, we propose an independent measurement of the observer’s peculiar velocity using strong gravitational lenses. Our method relies solely on angular measurements on the sky, providing a very robust way of extracting the observer velocity. The observable is derived purely from special relativity: a non-comoving observer perceives relativistic aberration, which induces a direction-dependent modulation of the observed Einstein radius. We demonstrate that selection effects, that modulate the number of detected Einstein rings across the sky, have a completely negligible impact on the measured dipole, below 0.1%. As such, this observable is insensitive to flux-related and clustering systematics that may bias source-count dipole experiments. This measurement, however, requires a very large number of strong gravitational lenses to detect a small distortion of the Einstein ring for lenses aligned with the observer velocity.
@@ 2 Einstein angle aberration
The peculiar velocity of the observer, $\bm{v}_{o}$, leads to a distortion of the Einstein ring due to aberration, as described in [Sereno2008, Dalang2023]. Consequently, a perfectly circular Einstein ring of radius $\theta_{E}$ will appear to a boosted observer as an ellipse, as shown in Fig. 1. To compute the distortion, without loss of generality, we align the $\bm{e}_{z}$ axis along the observer velocity and we place the direction of the center of mass of the lens $\bm{\hat{n}}$ in the $\bm{e}_{z}-\bm{e}_{y}$ plane, as illustrated in Figure 1. A boosted observer will see the center of mass in the direction $\bm{\hat{n}}^{\prime}\neq\bm{\hat{n}}$, which however still lies in the $\bm{e}_{z}-\bm{e}_{y}$ plane. The Einstein ring seen by the boosted observer lives in the plane orthogonal to $\bm{\hat{n}}^{\prime}$. To describe its shape, we adopt the same coordinate system as in [Dalang2023], i.e. we define $\bm{\hat{e}}_{1}^{\prime}$ along the $\bm{e}_{x}$ direction, such that angles along that axis are preserved under the boost, and $\bm{\hat{e}}_{2}^{\prime}$ lies in the $\bm{e}_{z}-\bm{e}_{y}$ plane, orthogonally to $\bm{\hat{n}}^{\prime}$. Angles along $\bm{\hat{e}}_{2}^{\prime}$ are deformed by the boost. We denote by $\theta_{\rm cm}$ the angle between $\bm{v}_{o}$ and the unboosted center of mass of the lens, and by $\theta_{\rm min}$ and $\theta_{\rm max}$ the angle between $\bm{v}_{o}$ and the unboosted extremities of the ring. The impact of the boost on an angle $\theta$, at linear order in $v_{o}/c$, is given by
@@
With this, we find that the boosted Einstein radius in direction $\bm{\hat{e}}_{2}^{\prime}$ is given by
@@
The fact that the boosted ring remains centered on the boosted center of mass is valid in the limit $\theta_{E2}\ll\theta_{\rm cm}$. In contrast, in the $\bm{\hat{e}}_{1}^{\prime}$ direction the boosted Einstein radius is the same as the unboosted one
$\theta^{\prime}_{E1}=\theta_{E1}\,.$ (2.3)
@@
Figure 1: Left panel: Coordinate system used in this work. We consider two observers, located at the origin $\mathcal{O}$. The first observer is stationary in this frame and measures comoving quantities, which are written without primes, e.g. $\theta_{\rm cm}$. Because of the symmetry around the $\bm{e}_{z}$ axis, the system is fully captured by lenses which lie in the $\bm{e}_{y}-\bm{e}_{z}$ plane. The stationary observer sees a perfect circular Einstein ring with Einstein angle $\theta_{E}$ in the direction $\bm{\hat{n}}$. They associate a coordinate system orthogonal to $\bm{\hat{n}}$ on the sky $\{\bm{\hat{e}}_{1},\bm{\hat{e}}_{2}\}$, depicted in the middle panel. This coordinate system is such that the $\bm{\hat{e}}_{1}$ vector is orthogonal both to $\bm{\hat{n}}$ and $\bm{v}_{o}$ (here it is aligned with $\bm{e}_{x}$), while $\bm{\hat{e}}_{2}$ lives in the $\bm{e}_{y}-\bm{e}_{z}$ plane. The second observer moves with peculiar velocity $\bm{v}_{o}$ aligned with $\bm{\hat{e}}_{z}$ and observes the Einstein ring in the $\bm{\hat{n}}^{\prime}$ direction, which forms an angle $\theta^{\prime}_{\rm cm}$ with the $\bm{e}_{z}$ axis. For this moving observer, the Einstein ring appears squashed along the $\bm{\hat{e}}^{\prime}_{2}$ direction, but unchanged along the $\bm{\hat{e}}^{\prime}_{1}$ direction. It appears approximately as an ellipse in the coordinate system defined by $\{\bm{\hat{e}}_{1}^{\prime},\bm{\hat{e}}_{2}^{\prime}\}$, as can be seen in the right panel. The Einstein angle as a function of the observed boosted azimuthal angle $\delta^{\prime}$ can be computed from the components $\theta^{\prime}_{E1}(\delta^{\prime})$ and $\theta^{\prime}_{E2}(\delta^{\prime})$. The amplitude of this deformation is proportional to $\cos(\theta_{\rm cm})v_{o}/c$.
The Einstein angle in an arbitrary direction can be expressed as a function of the azimuthal angle $\delta^{\prime}$, defined relative to $\bm{\hat{e}}^{\prime}_{1}$ :
$\theta^{\prime}_{E}(\delta^{\prime})$ $=\sqrt{\theta^{\prime}_{E1}(\delta^{\prime})^{2}+\theta^{\prime}_{E2}(\delta^{\prime})^{2}}$
$=\sqrt{\theta_{E1}(\delta^{\prime})^{2}+\theta_{E2}(\delta^{\prime})^{2}\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)^{2}}.$ (2.4)
For simplicity, we assume that, for an observer at rest, the Einstein ring is a perfect circle, arising from a spherical mass distribution, such as a Singular Isothermal Sphere (SIS)
$\theta_{E1}(\delta^{\prime})$ $=\theta_{E}\cos\delta^{\prime}\,$, (2.5)
$\theta_{E2}(\delta^{\prime})$ $=\theta_{E}\sin\delta^{\prime}\,.$ (2.6)
Inserting (2.5) and (2.6) into Eq. (2), we can relate the observed Einstein angle $\theta^{\prime}_{E}(\delta^{\prime})$ seen by a moving observer to the comoving one $\theta_{E}$
$\theta^{\prime}_{E}(\delta^{\prime})$ $=\theta_{E}\left(1-\sin^{2}\delta^{\prime}\,\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)\,.$ (2.7)
In practice, the Einstein angle is measured by fitting a mass model, e.g. a SIS model, to the imaging data. Consequently, the measured Einstein angle is azimuthally average over $\delta^{\prime}$:
$\theta^{\prime}_{E}$ $=\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{d}\delta^{\prime}\;\theta_{E}\left(1-\sin^{2}\delta^{\prime}\,\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$
$=\theta_{E}\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)\,.$ (2.8)
In the rest of this work, we use the average Einstein angle $\theta^{\prime}_{E}$ as our primary observable to measure $\bm{v}_{o}$.
@@
To increase the statistics, we measure the mean size of all detected rings in a given solid angle $\mathrm{d}\Omega^{\prime}$. Since the Einstein angle depends on the redshifts of the sources, $z^{\prime}_{s}$, and of the lens, $z^{\prime}_{l}$, this introduces two integrals weighted by the number of systems observed in that solid angle $\frac{\mathrm{d}N(\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})}{\mathrm{d}\Omega^{\prime}\mathrm{d}z^{\prime}_{s}\mathrm{d}z^{\prime}_{l}}$
@@
To evaluate Eq. (2.9), we need to express the Einstein radius as a function of the boosted redshifts $z^{\prime}_{s}$ and $z^{\prime}_{l}$, which introduces additional distortions. The Einstein radius depends indeed on the distances to the source and to the lens, and fixed values of $z^{\prime}_{s}$ and $z^{\prime}_{l}$ correspond to different distances in different directions. As shown in Appendix A, the Einstein radius in redshift space is given by
$\theta^{\prime}_{E}(\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=$ $\theta_{E}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)+\left((1+z_{s})\frac{\partial\theta_{E}}{\partial z_{s}}+(1+z_{l})\frac{\partial\theta_{E}}{\partial z_{l}}\right)\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\,$, (2.11)
where $\theta_{E}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))$ is the unboosted Einstein ring evaluated at the comoving distances $r(z^{\prime}_{s})$ and $r(z^{\prime}_{l})$, which do not depend on direction. [^1] Note that in the second term, we can use interchangeably $z$ or $z^{\prime}$ since they multiply a term which is already linear in $v_{o}/c$ From Eq. (2.11) we see that if we would measure the Einstein ring in bins of observed redshifts, the dipolar contribution would be affected by two terms: aberration and distance perturbations. However, once we integrate over all redshifts, the second contribution cancels and only aberration persists. To see this we need to express the number of detected Einstein rings in terms of $z_{s}^{\prime}$ and $z^{\prime}_{l}$. The detailed calculation can be found in Appendix A. We obtain
@@
The second line in Eq. (2.12) combines with the $2v_{o}/c$ in the first line, and with the second term in Eq. (2.11) to give rise to a boundary term, which exactly vanishes (see Appendix A for more detail). Moreover the effect of aberration in the number of detected systems (the remaining $2v_{o}/c$ in the first line of Eq. (2.12)) cancels in the numerator and the denominator of Eq. (2.9). Due to aberration, we see more Einstein rings per solid angle in the direction of the boost, but since we normalise the Einstein radius by the number of detected Einstein rings, this effect does not impact the final result. With this, we obtain
$\langle\theta^{\prime}_{E}(\bm{\hat{n}}^{\prime})\rangle=\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)\langle\theta_{E}\rangle\,$, (2.13)
@@
Eq. (2.13) allows us to measure the observer velocity $v_{o}$ by fitting a monopole term, proportional to the mean of the lens sample $\langle\theta_{E}\rangle$ and a dipole term proportional to $v_{o}$. This relies on the underlying assumptions that there are no selection effects, i.e. that all systems are detected. In Section 5, we will relax this assumption, introducing a selection in flux and in size. We will show that this introduces additional contributions to the dipole, which are however much smaller than the effect of aberration. More precisely the selection in size introduces a contamination of the order of 0.1% while the selection in flux is three orders of magnitude smaller. This shows that Einstein rings provide a very robust observable to measure the observer velocity, without being contaminated by selection effects that are typically difficult to characterize.
@@ 3.1 Likelihood definition
$\ln\mathcal{L}_{1}=-\frac{1}{2}\sum_{n=1}^{N}\frac{\left[\theta^{\prime}_{E_{n}}-\langle\theta_{E}\rangle\left(1-\frac{v_{o}}{2c}\cos(\theta^{\prime}_{{\rm cm}_{,n}})\right)\right]^{2}}{\delta\theta^{{}^{\prime}2}_{E_{n}}+\sigma_{\theta_{E}}^{2}}-\frac{1}{2}\sum_{n=1}^{N}\ln\!\big[2\pi\big(\delta\theta^{{}^{\prime}2}_{E_{n}}+\sigma_{\theta_{E}}^{2}\big)\big]\,$, (3.1)

### 19 Mar 2026 v2

@@ Abstract
Measurements of the kinematic cosmic dipole continue to show an intriguing tension between the value inferred from the CMB and that obtained from high-redshift source number counts. While the measured dipole direction appears consistent, the amplitude, set by the observer’s peculiar velocity $v_{o}$, remains in significant disagreement. In this paper, we propose using strong gravitational lenses with well-measured Einstein radii to estimate the kinematic cosmic dipole, through the relativistic aberration of the Einstein angle induced by the observer’s motion. We show that this effect could be detected solely from measurements of the Einstein radius in wide, high-resolution imaging surveys such as Euclid. However, the precision achievable using Einstein-radius measurements alone, without redshift or lens-galaxy mass information, appears insufficient to discriminate between the CMB value of $v_{o}$ and that derived from source number counts at high statistical significance. Nevertheless, we demonstrate that including a large sample of lenses with available kinematic information, either via the Fundamental Plane relation or, ideally, through spectroscopic velocity-dispersion measurements, drastically reduces the noise and substantially improves the constraining power of this method. We show that, for a realistic sample of strong lenses detected by Euclid and complemented with spectroscopic velocity dispersion measurements from 4MOST or DESI, it is possible to discriminate between the CMB- and source-number-counts-inferred values at the $\sim 5\sigma$ level using a new, fully independent method. We further demonstrate that this technique is only weakly sensitive to strong-lensing selection effects, with selection biases and threshold effects estimated to be well below the 1% level.
@@ 1 Introduction
In this work, we propose an independent measurement of the observer’s peculiar velocity using strong gravitational lenses. Our method relies solely on angular measurements on the sky, providing a very robust way of extracting the observer velocity. The observable is derived purely from special relativity: a non-comoving observer perceives relativistic aberration, which induces a direction-dependent modulation of the observed Einstein radius. We demonstrate that selection effects, which modulate the number of detected Einstein rings across the sky, have a completely negligible impact on the measured dipole, below 0.1%. As such, this observable is insensitive to flux-related and clustering systematics that may bias source-count dipole experiments. This measurement, however, requires a very large number of strong gravitational lenses to detect a small change in the Einstein ring’s size for lenses aligned with the observer velocity.
@@ 2 Einstein angle aberration
The peculiar velocity of the observer, $\bm{v}_{o}$, leads to a change in the size of the Einstein ring due to aberration, as described in [Sereno2008, Dalang2023]. Without loss of generality, we align the $\bm{e}_{z}$ axis along the observer’s velocity, and we place the direction of the center of mass of the lens $\bm{\hat{n}}$ in the $\bm{e}_{z}-\bm{e}_{y}$ plane, as illustrated in Figure 1. A boosted observer will see the center of mass in the direction $\bm{\hat{n}}^{\prime}\neq\bm{\hat{n}}$, which, however, still lies in the $\bm{e}_{z}-\bm{e}_{y}$ plane. The Einstein ring seen by the boosted observer lies in the plane orthogonal to $\bm{\hat{n}}^{\prime}$. To describe its shape, we adopt the same coordinate system as in [Dalang2023], i.e., we define $\bm{\hat{e}}_{1}^{\prime}$ along the $\bm{e}_{x}$ direction, such that angles along that axis are preserved under the boost, and $\bm{\hat{e}}_{2}^{\prime}$ lies in the $\bm{e}_{z}-\bm{e}_{y}$ plane, orthogonally to $\bm{\hat{n}}^{\prime}$. Angles along $\bm{\hat{e}}_{2}^{\prime}$ are deformed by the boost. We denote by $\theta_{\rm cm}$ the angle between $\bm{v}_{o}$ and the unboosted center of mass of the lens, and by $\theta_{\rm min}$ and $\theta_{\rm max}$ the angle between $\bm{v}_{o}$ and the unboosted extremities of the ring. The impact of the boost on an angle $\theta$, at linear order in $v_{o}/c$, is given by
@@
With this, we find that the boosted Einstein radius in the direction $\bm{\hat{e}}_{2}^{\prime}$ is given by
@@
The fact that the boosted ring remains centered on the boosted center of mass is valid in the limit $\theta_{E2}\ll\theta_{\rm cm}$. The boosted Einstein radius in direction $\bm{\hat{e}}_{1}^{\prime}$ is given by
$\theta^{\prime}_{E1}=\sin\theta^{\prime}_{\rm cm}(\varphi_{\rm max}-\varphi_{\rm cm})=\sin\theta^{\prime}_{\rm cm}(\varphi_{\rm cm}-\varphi_{\rm min})\,$, (2.3)
where $\varphi$ denotes the azimuthal angles of the center of mass and of the two extremities of the ring, which are unaffected by the boost. In the first equality, we have, as before, used that the size of the ring is much smaller than $\theta_{\rm cm}$. In our coordinate system, $\varphi_{\rm cm}=0$ and $\varphi_{\rm max}=\varphi=-\varphi_{\rm min}$. From Eq. (2.3), we see that even though angles along $\bm{\hat{e}}_{1}^{\prime}$ are unaffected by the boost, the Einstein radius along that direction is impacted due to the fact that $\theta^{\prime}_{\rm cm}$ changes under the boost. This may be understood visually from the second panel of Fig. 1. Inserting (2.1) into (2.3) and expanding at linear order in $v_{o}/c$, we find
$\theta^{\prime}_{E1}=\theta_{E1}\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)\,.$ (2.4)
The boost therefore produces an isotropic distortion of the ring [^1] At second order, and thus not relevant for this work, the transverse velocity between the lens and the observer can distort a perfectly circular Einstein ring into an ellipse, as shown in [Sereno2008]. In the following, we assume that the Einstein ring is a perfect circle, arising from a spherical mass distribution, such as a Singular Isothermal Sphere (SIS), while the introduction of ellipticity in the mass profile would simply increase the noise of our measurement (see Section 5). Since the distortion is isotropic, after the boost the ring remains a circle, but with a different radius
$\theta^{\prime}_{E}=\theta_{E}\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)\,.$ (2.5)
The boosted Einstein radius, $\theta^{\prime}_{E}$, is the primary observable in this work. In practice, this quantity is inferred by fitting a mass model to the imaging data and computing the azimuthally averaged Einstein radius of the resulting mass profile.
@@
Figure 1: Left panel: Coordinate system used in this work. We consider two observers, located at the origin $\mathcal{O}$. The first observer is stationary in this frame and measures comoving quantities, which are written without primes, e.g. $\theta_{\rm cm}$. Because of the symmetry around the $\bm{e}_{z}$ axis, the system is well captured by lenses which lie in the $\bm{e}_{y}-\bm{e}_{z}$ plane. The stationary observer sees a perfect circular Einstein ring with Einstein angle $\theta_{E}$ in the direction $\bm{\hat{n}}$. They associate a coordinate system orthogonal to $\bm{\hat{n}}$ on the sky $\{\bm{\hat{e}}_{1},\bm{\hat{e}}_{2}\}$, depicted in the third panel. This coordinate system is such that the $\bm{\hat{e}}_{1}$ vector is orthogonal both to $\bm{\hat{n}}$ and $\bm{v}_{o}$ (here it is aligned with $\bm{e}_{x}$), while $\bm{\hat{e}}_{2}$ lives in the $\bm{e}_{y}-\bm{e}_{z}$ plane. While angles along $\varphi$ are not affected by the boost ($\Delta\varphi\equiv\varphi_{\rm max}-\varphi_{\rm min}=\Delta\varphi^{\prime}$), the Einstein ring along $\bm{\hat{e}}_{1}$ is. This is due to the fact that $\theta_{\rm cm}^{\prime}$ is affected by the boost, as may be understood from the second panel. The second observer moves with peculiar velocity $\bm{v}_{o}$ aligned with $\bm{\hat{e}}_{z}$ and observes the Einstein ring in the $\bm{\hat{n}}^{\prime}$ direction, which forms an angle $\theta^{\prime}_{\rm cm}$ with the $\bm{e}_{z}$ axis. For this moving observer, the Einstein ring appears smaller, as may be seen in the fourth panel. This deformation preserves the shape of the Einstein ring and is proportional to $\cos(\theta_{\rm cm})v_{o}/c$.
@@
To increase the statistics, we measure the mean size of all detected rings in a given solid angle $\mathrm{d}\Omega^{\prime}$. Since the Einstein angle depends on the redshifts of the sources $z^{\prime}_{s}$ and of the lens $z^{\prime}_{l}$, this introduces two integrals weighted by the number of systems observed in that solid angle $\frac{\mathrm{d}N(\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})}{\mathrm{d}\Omega^{\prime}\mathrm{d}z^{\prime}_{s}\mathrm{d}z^{\prime}_{l}}$
@@
To evaluate Eq. (2.6), we need to express the Einstein radius as a function of the boosted redshifts $z^{\prime}_{s}$ and $z^{\prime}_{l}$, which introduces additional distortions. The Einstein radius indeed depends on the distances to the source and to the lens, and fixed values of $z^{\prime}_{s}$ and $z^{\prime}_{l}$ correspond to different distances in different directions. As shown in Appendix A, the Einstein radius in redshift space is given by
$\theta^{\prime}_{E}(\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=$ $\theta_{E}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)+\left((1+z_{s})\frac{\partial\theta_{E}}{\partial z_{s}}+(1+z_{l})\frac{\partial\theta_{E}}{\partial z_{l}}\right)\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\,$, (2.8)
where $\theta_{E}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))$ is the unboosted Einstein ring evaluated at the comoving distances $r(z^{\prime}_{s})$ and $r(z^{\prime}_{l})$, which do not depend on direction [^2] Note that in the second term, we can use $z$ or $z^{\prime}$ interchangeably since they multiply a term that is already linear in $v_{o}/c$. From Eq. (2.8), we see that if we were to measure the Einstein ring in bins of observed redshifts, the dipolar contribution would be affected by two terms: aberration and distance perturbations. However, once we integrate over all redshifts, the second contribution cancels, and only aberration persists. To see this, we need to express the number of detected Einstein rings in terms of $z_{s}^{\prime}$ and $z^{\prime}_{l}$. The detailed calculation can be found in Appendix A. We obtain
@@
The second line in Eq. (2.9) combines with the $2v_{o}/c$ in the first line, and with the second term in Eq. (2.8) to give rise to a boundary term that exactly vanishes (see Appendix A for more detail). Moreover, the effect of aberration in the number of detected systems (the remaining $2v_{o}/c$ in the first line of Eq. (2.9)) cancels in the numerator and the denominator of Eq. (2.6). Due to aberration, we see more Einstein rings per solid angle in the direction of the boost, but since we normalize the Einstein radius by the number of detected Einstein rings, this effect does not impact the final result. With this, we obtain
$\langle\theta^{\prime}_{E}(\bm{\hat{n}}^{\prime})\rangle=\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)\langle\theta_{E}\rangle\,$, (2.10)
@@
Eq. (2.10) allows us to measure the observer velocity $v_{o}$ by fitting a monopole term, proportional to the mean of the lens sample $\langle\theta_{E}\rangle$, and a dipole term proportional to $v_{o}$. This relies on the underlying assumptions that there are no selection effects, i.e. that all systems are detected. In Section 5, we will relax this assumption, introducing a selection in flux and in size. We will show that this introduces additional contributions to the dipole, which are, however, much smaller than the effect of aberration. More precisely, the selection in size introduces a contamination of the order of 0.1%, while the selection in flux is three orders of magnitude smaller. This shows that Einstein rings provide a very robust observable to measure the observer velocity, without being contaminated by selection effects that are typically difficult to characterize.
@@ 3.1 Likelihood definition
$\ln\mathcal{L}_{1}=-\frac{1}{2}\sum_{n=1}^{N}\frac{\left[\theta^{\prime}_{E_{n}}-\langle\theta_{E}\rangle\left(1-\frac{v_{o}}{c}\cos(\theta^{\prime}_{{\rm cm}_{,n}})\right)\right]^{2}}{\delta\theta^{{}^{\prime}2}_{E_{n}}+\sigma_{\theta_{E}}^{2}}-\frac{1}{2}\sum_{n=1}^{N}\ln\!\big[2\pi\big(\delta\theta^{{}^{\prime}2}_{E_{n}}+\sigma_{\theta_{E}}^{2}\big)\big]\,$, (3.1)